When heated, lithium reacts with nitrogen to form lithium nitride: (4 points) Li(s) + N2(g) ----------> Li3N(s) What is the theoretical yield of Li3N in grams when 12.3 g of Li are heated with 33.6 g of N2? If the actual yield of Li3N is 5.89 g, what is the percent yield of the reaction?


1
Expert's answer
2021-02-04T07:04:25-0500

6Li + N2 = 2Li3N

"\nu" (Li)=m/M=12.3/7=1.76 mol

"\nu"(N2)=33.6/28= 1.2 mol

"\nu"(Li3N)= 1.76*2/6=0.59 mol

m(Li3N)= 0.59*35=20.65 g

"\eta" = mp/mt*100%=5.89/20.65*100%=29%